# Investigating the correlation between ZTF TDEs and IceCube high-energy neutrinos

### v2

@@ Abstract
The main contributors of the IceCube diffuse neutrino flux remain unclear. Tidal disruption events (TDEs) have been proposed as potential emitters of the high-energy neutrinos detected by IceCube. Therefore, investigating the correlation between the TDE population and IceCube neutrinos could help us better understand whether the TDE population could be potential high-energy neutrino emitters. In this paper, we perform a systematic search for TDEs that are associated with neutrinos in a sample including 143 IceCube neutrino alert events and 56 TDEs classified by the Zwicky Transient Facility (ZTF) - Bright Transient Survey (BTS). Furthermore, considering that the TDEs/TDE candidates reported as potential IceCube neutrino emitters are all accompanied by infrared (IR) observations, we further select the TDEs with IR observations from these 56 TDEs as a subsample to examine the correlation with neutrinos. Based on the Wide-field Infrared Survey Explorer (WISE) mission database, eight TDEs are identified as having IR observations. Since good spatial localization is crucial for association analysis, we employ two methods to handle alert events with large error radii in our sample. Then we employ Monte Carlo simulations to quantify the correlation between the TDE sample/subsample and IceCube neutrinos. Finally, after considering spatial and temporal criteria, the eight TDEs with IR echoes show the most significant correlation at a 2.37 confidence level. If we tentatively further take the time delay factor into account in the weighting scheme, the correlation enhances to 2.65 confidence level.
@@ 1 Introduction
The detection of the TeV-PeV diffuse astrophysical neutrino flux by IceCube (Aartsen et al., 2013a, 2014a, 2015, 2016, 2020a; Abbasi et al., 2021a, 2022a) has unveiled new possibilities for exploration in both astrophysics (Kistler & Beacom, 2006; Beacom & Kistler, 2007; Murase et al., 2011; Murase & Ioka, 2013; Murase et al., 2013; Ahlers & Halzen, 2014; Tamborra et al., 2014; Murase et al., 2014; Bechtol et al., 2017; Kistler & Laha, 2018; Bartos et al., 2017; Sudoh et al., 2018; Bartos et al., 2019; Bustamante & Ahlers, 2019; Hovatta et al., 2021; Ackermann et al., 2022) and particle physics (Gonzalez-Garcia et al., 2005; Ioka & Murase, 2014; Ng & Beacom, 2014; Aartsen et al., 2017a; Bustamante & Connolly, 2019; Zhou & Beacom, 2020a, b, 2022; Aartsen et al., 2021; Ackermann et al., 2022; Guo et al., 2023; Plestid & Zhou, 2024; Lü et al., 2024). Despite a significant number of studies have been conducted to identify the sources of these astrophysical neutrinos (Abbasi et al., 2011; Aartsen et al., 2013b, 2014b; Adrian-Martinez et al., 2016; Aartsen et al., 2017b, c, 2019a, 2020b, 2019b, 2020c; Lu et al., 2024; Kouch et al., 2024), it remains unclear which astrophysical sources are the primary origins of these neutrinos, and only a very small number of sources (e.g., the Seyfert galaxy NGC 1068) have been identified as neutrino sources (Abbasi et al., 2022b).
@@
Among the TDEs/TDE candidates that have been reported as potential IceCube neutrino emitters, three notable events include: one identified TDE AT2019dsg (Stein et al., 2021), and two TDE candidates, AT2019fdr (Reusch et al., 2022) and AT2019aalc (van Velzen et al., 2024), which correspond to the possibly associated alert events IC191001A, IC200530A, and IC19119A, respectively. In these three events, AT2019dsg has been classified as a member of TDE by the Zwicky Transient Facility (ZTF). The long-lived nonthermal emission detected in AT2019dsg (Cendes et al., 2021; Stein et al., 2021) suggests that mildly relativistic outflows may provide an ideal environment for neutrino production. In contrast, the classification of AT2019fdr as a TDE is somewhat uncertain, as the flare occurred in a known active galactic nucleus (AGN), complicating its origin due to the presence of an accretion disk (e.g., Trakhtenbrot et al., 2019; Frederick et al., 2021; Cendes et al., 2021). Similarly, AT2019aalc is also posited to be a flare originating from another known AGN (van Velzen et al., 2024; Veres et al., 2024). These three events share several comparable characteristics; for instance, they have bright optical luminosities, accompanying infrared (IR) echoes, i.e., delayed IR rising after the discovery date of TDE. Further, all of these events have a time delay for the neutrino detection, which has been explained by some models, for instance, protons generated by the delayed choked jet (Mukhopadhyay et al., 2024), the slow acceleration of protons by the plasma/magnetic reconnection (Murase et al., 2020), an increase of the neutrino production efficiency over time assuming that the radius of neutrino production evolves similarly to that of the black body (Winter & Lunardini, 2021), and the protons accelerated by the outflow/jet interacting with protons/photons after traveling to the cloud/dust region (Wu et al., 2022; Winter & Lunardini, 2023).
@@
Besides these three sources, there are other TDEs / TDE candidates that have also been reported as possible IceCube high-energy neutrino emitters. For instance, Jiang et al. (2023) reported two obscured TDE candidates as potential neutrino emitters based on a sample of mid-IR outbursts in nearby galaxies (MIRONG). Yuan et al. (2024) reported a super-bright TDE candidate with IR emission that satisfies spatial-temporal coincidence when considering systematic errors. Recently, Li et al. (2024) reported another TDE that is associated with a potential high-energy neutrino flare at a 2.9 confidence level, which also accompanies IR observation.
@@
As more and more TDEs/TDE candidates are reported as potential IceCube high-energy neutrino emitters, investigating the correlation between the TDE population and high-energy neutrinos is helpful for understanding whether the TDE population could indeed be contributors to the IceCube neutrino flux. In this work, we investigate the hypothesis that the TDE population is correlated with IceCube alert events111https://gcn.gsfc.nasa.gov/amon_icecube_gold_bronze_events.html. We utilize the TDEs classified by ZTF - Bright Transient Survey (BTS)222https://sites.astro.caltech.edu/ztf/bts/bts.php (Perley et al., 2020), which is a public catalog with a strict pipeline to conduct a classification for transient sources, and a magnitude-limited ( in either the or filter) survey for extragalactic transients in the ZTF public stream (Perley et al., 2020). We systematically search for spatial association events between TDEs of the ZTF-BTS catalog and IceCube alert events, while allowing a time delay between TDEs and the alert events. Since IR echoes are detected from all the TDEs/TDE candidates that are reported as potential IceCube neutrino emitters, we also systematically search for the IR emission at the coordinates of ZTF-BTS TDEs using the Wide-field Infrared Survey Explorer (WISE) mission database (Wright et al., 2010). Thus, we further adopt TDEs with IR observations as a subsample to investigate the correlation with IceCube high-energy neutrinos.
@@ 2.1 TDE sample
Then we search for the IR echoes from coordinates of ZTF-BTS TDEs using the full Near-Earth Object WISE Reactivation database444https://irsa.ipac.caltech.edu/cgi-bin/Gator/nph-dd?catalog=neowiser_p1bs_psd (NEOWISE-R; Mainzer et al., 2014; NEOWISE Team, 2020). The full NEOWISE-R database is up to July 2024. To ensure that the IR echoes of TDEs can be included in the NEOWISE-R database, we limit our TDE sample to before May 2024, due to the fact that the IR echoes of TDEs often have significant time delay. After the time cut, there are 56 TDEs remaining in our sample, as shown in Fig. 1. The NEOWISE-R database is based on an imaging survey centered at 3.4 and 4.6 m (labeled W1 and W2). The photometric measurements are obtained from the fitting of the point-spread function (PSF), where the PSFs are estimated from observations involving tens of thousands of stars. The NEOWISE-R magnitudes are measured through PSF fitting for individual exposures, which have been utilized to examine the IR flares for astrophysical sources (e.g., Jiang et al., 2021).
@@
where, and are the maximum and minimum magnitude for -th band, respectively. is the error of and is the error of . To identify the IR variability, we impose a requirement that the IR variability, denoted as or , must exceed 5. We used the following criteria to further select TDEs with IR echoes from these TDEs with IR variability. 1) The IR magnitude presents brightening compared to the median value of the quiescent level before the discovery date (Jiang et al., 2021). 2) The brightest magnitude should occur after the discovery dates of TDEs that are obtained from the Transient Name Server (TNS)555https://www.wis-tns.org/search. Finally, there are 8 ZTF-BTS TDEs with IR echoes, as listed in Table 1. The light curves of these eight TDEs are presented in Fig 2, where red and blue circles are the W1 and W2 band data, respectively. The red dashed lines are the discovery dates of TDEs.
@@ 3 Analysis Method
Based on these two methods for handling the alert sample and calculating the TS values, we perform Monte Carlo simulations to quantify the correlation between the TDE population and the alert population. Since in our previous work we found that different randomization methods in the MC simulations may have some effect on the results of the correlation analysis (Lu et al., 2024), in the following analysis we will test three different randomization methods.
@@
Method II: We not only randomize the RA of alert events, as in Method I, but also randomize the positions of the TDEs using the method described in Lu et al. (2024). For each mock TDE source, we randomly select an and a from the observed lists of RA and DEC, then impose an additional random shift to within a region of radius to obtain the position of the mock source. In the simulations, we require the simulated sources not to be outside the sky region of the ZTF-BTS sample, which covers the northern sky with a DEC range from to 90∘ but excluding the Galactic plane region (). We also require the number of mock sources to be the same as the real sample. Finally, we perform KS tests on and distributions to ensure that the mock sources preserve the same distribution pattern of the actual sources.
@@ 4 Results
We first systematically search for associated events that satisfy spatial and temporal criteria. After adopting the CUT method, there are 122 alert events remaining in our sample. Two associated events are found, and all of these events have IR echoes, which are listed in Table 2. Apart from AT2019dsg, the other TDE is associated with one bronze alert events.
Under the WEIGHT method, all 143 alert events are utilized to search for associated events. We found that there are six occurrences of TDEs within the of five alert events. Among them, one TDE (AT2019azh) is located within of two alert events. Additionally, the of two alert events each includes two TDEs. However, we note that the values of these two alert events are significantly large ( 10∘), contributing only negligible TS values. In total, five distinct TDEs are found within of four alert events, two of which have IR echoes, as listed in Table 3.
RAalert | DECalert | Signalness | TDE event | TS1 valueb | IR echoesc | ||||
(deg) | (deg) | (deg) | (days) | ||||||
124.54 | 20.74 | 2.39 | 23/02/17 | 0.454 | AT2019azh | 1456 | 1 | Yes | 0.022 |
314.08 | 12.94 | 2.95 | 19/10/01 | 0.589 | AT2019dsg | 175 | 1 | Yes | 0.0512 |
@@
RAalert | DECalert | Datealert | Signalness | TDE event | TS1 valuec | IR echoese | ||||
(deg) | (deg) | (deg) | (days) | |||||||
239.63 | 39.94 | 15.38 | 24/03/07 | 0.606 | AT2020vwl | 1244 | 14.19 | 7.90 | No | 0.0325 |
239.63 | 39.94 | 15.38 | 24/03/07 | 0.606 | AT2020pj | 1526 | 8.72 | 2.68 | No | 0.068 |
127.18 | 20.74 | 10.16 | 23/07/07 | 0.466 | AT2022bdw | 522 | 2.31 | 8.78 | No | 0.03782 |
127.18 | 20.74 | 10.16 | 23/07/07 | 0.466 | AT2019azh | 1596 | 4.06 | 6.85 | Yes | 0.022 |
124.54 | 20.74 | 2.39 | 23/02/17 | 0.454 | AT2019azh | 1456 | 2.22 | 2.39 | Yes | 0.022 |
314.08 | 12.94 | 2.95 | 19/10/01 | 0.589 | AT2019dsg | 175 | 1.28 | 9.63 | Yes | 0.0512 |
@@
CUT method | 0.39 | 0.38 | 0.37 | All TDEs |
WEIGHT method | 0.40 | 0.42 | 0.42 | All TDEs |
CUT method | 0.025 | 0.018 | 0.022 | TDEs with IR echoes |
WEIGHT method | 0.038 | 0.044 | 0.055 | TDEs with IR echoes |
By realizing 104 times MC simulations, the results obtained from the three MC methods are summarized in Table 4, which includes the results from both the CUT method and the WEIGHT method. Using the all TDEs sample, the best -values are 0.37 and 0.40 for the CUT method and WEIGHT method, respectively. When using the subsample of TDEs with IR echoes, the best -values are 0.018 and 0.038 for the CUT method and WEIGHT method, corresponding to 2.37 and 2.08, respectively. The most significant correlation with neutrinos is obtained from the subsample of TDEs with IR echoes at a confidence level, which may indicate that TDEs with IR echoes are more likely to correlate with IceCube neutrinos compared to all TDEs sample.
5 Discussion

### v3

@@ Abstract
Investigating the correlation between the TDE population and IceCube neutrinos could help us better understand whether TDEs could be potential high-energy neutrino emitters. In this paper, we perform a systematic search for TDEs that are associated with neutrinos in a sample including 143 IceCube neutrino alert events and 52 TDEs classified by the Zwicky Transient Facility (ZTF) - Bright Transient Survey (BTS). Furthermore, considering that the TDEs/TDE candidates reported as potential IceCube neutrino emitters are all accompanied by infrared (IR) echo emissions, we further select the TDEs with IR echoes from these 52 TDEs as a subsample to examine the correlation with neutrinos. Based on the Wide-field Infrared Survey Explorer (WISE) mission database, seven TDEs are identified as having IR echoes. Then we employ Monte Carlo simulations to quantify the correlation between the TDE sample/subsample and IceCube neutrinos. Finally, after considering spatial and temporal criteria, the seven TDEs with IR echoes show the most significant correlation at a 2.46 confidence level. If we tentatively further take the time delay factor into account, the correlation enhances to a 2.66 confidence level. The correlation is primarily contributed by two TDEs: AT2019dsg and AT2019azh. The latter’s association with a neutrino alert, IC230217A, is newly reported in this work. We discussed the possible physical connection between AT2019azh and the neutrino event IC230217A.
@@ 1 Introduction
The detection of the TeV-PeV diffuse astrophysical neutrino flux by IceCube (Aartsen et al., 2013a, 2014a, 2015, 2016, 2020a; Abbasi et al., 2021a, 2022a) has unveiled new possibilities for exploration in both astrophysics (Kistler & Beacom, 2006; Beacom & Kistler, 2007; Murase et al., 2011; Murase & Ioka, 2013; Murase et al., 2013; Ahlers & Halzen, 2014; Tamborra et al., 2014; Murase et al., 2014; Bechtol et al., 2017; Kistler & Laha, 2018; Bartos et al., 2017; Sudoh et al., 2018; Bartos et al., 2019; Bustamante & Ahlers, 2019; Hovatta et al., 2021; Ackermann et al., 2022) and particle physics (Gonzalez-Garcia et al., 2005; Ioka & Murase, 2014; Ng & Beacom, 2014; Aartsen et al., 2017a; Bustamante & Connolly, 2019; Zhou & Beacom, 2020a, b, 2022; Aartsen et al., 2021; Ackermann et al., 2022; Guo et al., 2023; Plestid & Zhou, 2024; Lü et al., 2024). Despite a significant number of studies have been conducted to identify the sources of these astrophysical neutrinos (Abbasi et al., 2011; Aartsen et al., 2013b, 2014b; Adrian-Martinez et al., 2016a; Aartsen et al., 2017b, c, 2019a, 2020b, 2019b, 2020c; Zhou et al., 2021; Chang et al., 2024; Kouch et al., 2024), it remains unclear which astrophysical sources are the primary origins of these neutrinos, and only a very small number of sources (e.g., the Seyfert galaxy NGC 1068) have been identified as neutrino sources (Abbasi et al., 2022b).
@@
Among the TDEs/TDE candidates that have been reported as potential IceCube neutrino emitters, three notable events include: one identified TDE AT2019dsg (Stein et al., 2021), and two TDE candidates, AT2019fdr (Reusch et al., 2022) and AT2019aalc (van Velzen et al., 2024), which correspond to the possibly associated alert events IC191001A, IC200530A, and IC191119A, respectively. In these three events, AT2019dsg has been classified as a member of TDE by the Zwicky Transient Facility (ZTF). The long-lived nonthermal emission detected in AT2019dsg (Cendes et al., 2021; Stein et al., 2021) suggests that mildly relativistic outflows may provide an ideal environment for neutrino production. In contrast, the classification of AT2019fdr as a TDE is somewhat uncertain, as the flare occurred in a known active galactic nucleus (AGN), complicating its origin due to the presence of an accretion disk (e.g., Trakhtenbrot et al., 2019; Frederick et al., 2021; Cendes et al., 2021). Similarly, AT2019aalc is also posited to be a flare originating from another known AGN (van Velzen et al., 2024; Veres et al., 2024). These three events share several comparable characteristics; for instance, they have bright optical luminosities, accompanying infrared (IR) echoes, i.e., delayed IR rising after the discovery date of TDE. Further, all of these events have a time delay for the neutrino detection, which has been explained by some models, for instance, protons generated by the delayed choked jet precession (Mukhopadhyay et al., 2024; Lu et al., 2024b), the slow acceleration of protons by the plasma/magnetic reconnection (Murase et al., 2020), an increase of the neutrino production efficiency over time assuming that the radius of neutrino production evolves similarly to that of the black body (Winter & Lunardini, 2021), and the protons accelerated by the outflow/jet interacting with protons/photons after traveling to the cloud/dust region (Wu et al., 2022; Winter & Lunardini, 2023).
@@
Besides these three sources, there are other TDEs / TDE candidates that have also been reported as possible IceCube high-energy neutrino emitters. For instance, Jiang et al. (2023) reported two obscured TDE candidates as potential neutrino emitters based on a sample of mid-IR outbursts in nearby galaxies (MIRONG). Yuan et al. (2024) reported a super-bright TDE candidate with IR emission that satisfies spatial-temporal coincidence when considering systematic errors. Recently, Li et al. (2024) reported another TDE that is associated with a potential high-energy neutrino flare at a 2.9 confidence level, which also accompanies IR echoes.
@@
As more and more TDEs/TDE candidates are reported as potential IceCube high-energy neutrino emitters, investigating the correlation between the TDE population and high-energy neutrinos is helpful for understanding whether the TDE population could indeed be contributors to the IceCube neutrino flux. In this work, we investigate the hypothesis that the TDE population is correlated with IceCube alert events111https://gcn.gsfc.nasa.gov/amon_icecube_gold_bronze_events.html. We utilize the TDEs classified by ZTF - Bright Transient Survey (BTS)222https://sites.astro.caltech.edu/ztf/bts/bts.php (Perley et al., 2020), which is a public catalog with a strict pipeline to conduct a classification for transient sources, and a magnitude-limited ( in either the or filter) survey for extragalactic transients in the ZTF public stream (Perley et al., 2020). We systematically search for spatial association events between TDEs of the ZTF-BTS catalog and IceCube alert events, while allowing a time delay between TDEs and the alert events. Since IR echoes are detected from all the TDEs/TDE candidates that are reported as potential IceCube neutrino emitters, we also systematically search for the IR emission at the coordinates of ZTF-BTS TDEs using the Wide-field Infrared Survey Explorer (WISE) mission database (Wright et al., 2010). Thus, we further adopt TDEs with IR echoes as a subsample to investigate the correlation with IceCube high-energy neutrinos.
@@ 2.1 TDE sample
Then we search for the IR echoes from coordinates of ZTF-BTS TDEs using the full Near-Earth Object WISE Reactivation database444https://irsa.ipac.caltech.edu/cgi-bin/Gator/nph-dd?catalog=neowiser_p1bs_psd (NEOWISE-R; Mainzer et al., 2014; NEOWISE Team, 2020). The full NEOWISE-R database is up to July 2024. Most parts of the sky are visited every six months, receiving approximately 10 observations during a 24 h period of each visit (Wright et al., 2010). To ensure that the IR echoes of TDEs can be included in the NEOWISE-R database, we limit our TDE sample to before January 2024 (i.e., six months before the end of the NEOWISE-R data), due to the fact that the IR echoes of TDEs often have significant time delay and considering the six-month survey cadence of NEOWISE. After the time cut, there are 52 TDEs remaining in our sample, as shown in Fig. 1. The NEOWISE-R database is based on an imaging survey centered at 3.4 and 4.6 m (labeled W1 and W2). The photometric measurements are obtained from the fitting of the point-spread function (PSF), where the PSFs are estimated from observations involving tens of thousands of stars. The NEOWISE-R magnitudes are measured through PSF fitting for individual exposures, which have been utilized to examine the IR flares for astrophysical sources (e.g., Jiang et al., 2021).
@@
where, and are the maximum and minimum magnitude for -th band, respectively. is the error of and is the error of . To identify the IR variability, we impose a requirement that the IR variability, denoted as or , must exceed 5. We used the following criteria to further select TDEs with IR echoes from these TDEs with IR variability. 1) The IR magnitude presents brightening compared to the median value of the quiescent level before the discovery date (Jiang et al., 2021). 2) The brightest magnitude should occur after the discovery dates of TDEs that are obtained from the Transient Name Server (TNS)555https://www.wis-tns.org/search. Finally, there are 7 ZTF-BTS TDEs with IR echoes, as listed in Table 1. The light curves of these seven TDEs are presented in Fig 2, where red and blue circles are the W1 and W2 band data, respectively. The red dashed lines are the discovery dates of TDEs.
@@ 3 Analysis Method
Based on these two methods for handling the alert sample and calculating the TS values, we perform Monte Carlo simulations to quantify the correlation between the TDE population and the alert population. Since in our previous work we found that different randomization methods in the MC simulations may have some effect on the results of the correlation analysis (Lu et al., 2024a), in the following analysis we will test three different randomization methods.
@@
Method II: We not only randomize the RA of alert events, as in Method I, but also randomize the positions of the TDEs using the method described in Lu et al. (2024a). For each mock TDE source, we randomly select an and a from the observed lists of RA and DEC, then impose an additional random shift to within a region of radius to obtain the position of the mock source. In the simulations, we require the simulated sources not to be outside the sky region of the ZTF-BTS sample, which covers the northern sky with a DEC range from to 90∘ but excluding the Galactic plane region (). We also require the number of mock sources to be the same as the real sample. Finally, we perform KS tests on and distributions to ensure that the mock sources preserve the same distribution pattern of the actual sources.
@@ 4 Results
4.1 Evaluate the correlation between ZTF TDEs and neutrino alerts
We first systematically search for associated events that satisfy spatial and temporal criteria. After adopting the CUT method, there are 122 alert events remaining in our sample. Two associated events are found, and all of these events have IR echoes, which are listed in Table 2. Apart from AT2019dsg, another TDE (AT2019azh) is also found to be associated with a neutrino alert (IC230217A). The association between AT2019azh and IC230217A has not been reported before and is newly discovered in this work. We will further discuss their connection in Sec. 5.2.
Under the WEIGHT method, all 143 alert events are utilized to search for associations. We find that there are six occurrences of TDEs within the of five alert events. Among them, one TDE (AT2019azh) is located within of two alert events. Additionally, the of two alert events each includes two TDEs. However, we note that the values of these two alert events are significantly large ( 10∘), contributing only negligible TS values. In total, five distinct TDEs are found within of four alert events, two of which have IR echoes, as listed in Table 3. Apart from AT2019dsg, the other TDE that contributes a relatively large TS in the WEIGHT method is still AT2019azh.
Alert | RAalert | DECalert | Signalness | TDE event | TS1 valueb | IR echoesc | ||||
(deg) | (deg) | (deg) | (days) | |||||||
IC230217A | 124.54 | 20.74 | 2.39 | 23/02/17 | 0.454 | AT2019azh | 1456 | 1 | Yes | 0.022 |
IC191001A | 314.08 | 12.94 | 2.95 | 19/10/01 | 0.589 | AT2019dsg | 175 | 1 | Yes | 0.0512 |
@@ 4.1 Evaluate the correlation between ZTF TDEs and neutrino alerts
Alert | RAalert | DECalert | Datealert | Signalness | TDE event | TS1 valuec | IR echoese | ||||
(deg) | (deg) | (deg) | (yy/mm/dd) | (days) | (deg) | ||||||
IC240307A | 239.63 | 39.94 | 15.38 | 24/03/07 | 0.606 | AT2020vwl | 1244 | 14.19 | 7.90 | No | 0.0325 |
IC240307A | 239.63 | 39.94 | 15.38 | 24/03/07 | 0.606 | AT2020pj | 1526 | 8.72 | 2.68 | No | 0.068 |
IC230707B | 127.18 | 20.74 | 10.16 | 23/07/07 | 0.466 | AT2022bdw | 522 | 2.31 | 8.78 | No | 0.03782 |
IC230707B | 127.18 | 20.74 | 10.16 | 23/07/07 | 0.466 | AT2019azh | 1596 | 4.06 | 6.85 | Yes | 0.022 |
IC230217A | 124.54 | 20.74 | 2.39 | 23/02/17 | 0.454 | AT2019azh | 1456 | 2.22 | 2.39 | Yes | 0.022 |
IC191001A | 314.08 | 12.94 | 2.95 | 19/10/01 | 0.589 | AT2019dsg | 175 | 1.28 | 9.63 | Yes | 0.0512 |
@@
CUT method | 0.38 | 0.36 | 0.36 | All TDEs |
WEIGHT method | 0.39 | 0.42 | 0.41 | All TDEs |
CUT method | 0.026 | 0.014 | 0.018 | TDEs with IR echoes |
WEIGHT method | 0.039 | 0.042 | 0.052 | TDEs with IR echoes |
By realizing 104 times MC simulations, the results obtained from the three MC methods are summarized in Table 4, which includes the results from both the CUT method and the WEIGHT method. Using the all TDEs sample, the best -values are 0.36 and 0.39 for the CUT method and WEIGHT method, respectively. When using the subsample of TDEs with IR echoes, the best -values are 0.014 and 0.039 for the CUT method and WEIGHT method, corresponding to 2.46 and 2.06, respectively. The most significant correlation with neutrinos is obtained from the subsample of TDEs with IR echoes at a confidence level, which may indicate that TDEs with IR echoes are more likely to correlate with IceCube neutrinos compared to the all-TDE sample.